OpenStudy (anonymous):

please help explain how to solve this on the unit circle. interval (0, 2pi) for (sinx -1)(2sin^2x-5 sinx+2)=0

OpenStudy (nory):

Try using the 0-product property.

OpenStudy (anonymous):

I'm not sure what that is

OpenStudy (nory):

Remember in algebra when there was an equation like (x-1)(x-2)=0?

OpenStudy (anonymous):

how do I solve it on the unit circle

OpenStudy (nory):

Not exactly. The property is that if there is a multiplication that equals 0, then one or more factors must equal 0.

OpenStudy (anonymous):

i dont remember that. can you please explain further

OpenStudy (nory):

So for (x-1)(x-2)=0, then either x = 1 or x=2. That's how I would solve it.

OpenStudy (anonymous):

would x= 2pi,x=pi/2, x=5pi/4 be right?

OpenStudy (nory):

Try plugging it into the equation.